OpenStudy (rose21):

Terrence answered 5 homework questions in ten minutes. How long will it take him to answer twenty-five questions

OpenStudy (yttrium):

Use this proportionality. \[\frac{ 5 }{ 10 } = \frac{ 25 }{ x }\] Just solve for x :)

terenzreignz (terenzreignz):

Finally... a question with my name on it :D What didn't you get @rose21

OpenStudy (rose21):

lol

OpenStudy (yttrium):

hahahaha LOL @terenzreignz

terenzreignz (terenzreignz):

Except for those 2 r's. Who need double r anyway? :3 @rose21 what was your answer?

OpenStudy (rose21):

let me check

terenzreignz (terenzreignz):

Taking ten minutes to answer five questions... that doesn't sound like me at all :3 XD

OpenStudy (yttrium):

@terenzreignz is humble :D HAHAHAHA

terenzreignz (terenzreignz):

that wasn't what I meant at all... but sure... :3

OpenStudy (yttrium):

hahaha, maybe it depends upon the question XD

terenzreignz (terenzreignz):

Okay @rose21 Listen up :D \[\frac{ 5 }{ 10 } = \frac{ 25 }{ x }\] To solve this, you cross-multiply. Left numerator times right-denominator is equal to left-denominator times right-numerator. got it?